Determine the Taylor Series for the function f(x) = e-3 centered at α = -1. ΑΣ-3)* * (t+ 1): Β. Σ (a + 1): «Σ (a + 1) b.
Answer #1

We know no we are asked to find the Taylon deries of for) centered at / f(x) = x - 30 can be written as f(x) = &- 3(x+1-1) (-n=0 é oe! - 1 t t (option o)